OpenStudy (anonymous):

What is the first step in solving the equation three-tenths equals x divided by three plus one-tenth for x? a. Add one-tenth to both sides of the equal sign. b. Add three-tenths to both sides of the equal sign. c. Subtract one-tenth from both sides of the equal sign. d. Subtract three-tenths from both sides of the equal sign.

OpenStudy (whpalmer4):

\[\frac{3}{10} = \frac{x}{3}+\frac{1}{10}\] You're trying to make the equation simpler. Do any of those choices do that?

OpenStudy (anonymous):

would you subtract \[\frac{ 1 }{ 10 }\] from both sides

OpenStudy (whpalmer4):

Exactly! Then you would have \[\frac{3}{10}-\frac{1}{10} = \frac{x}{3}+\frac{1}{10}-\frac{1}{10}\]\[\frac{2}{10}=\frac{x}{3}\]

OpenStudy (whpalmer4):

Cross multiply and solve for \(x\) would be the next step

OpenStudy (whpalmer4):

\[2*3 = 10*x\]

OpenStudy (whpalmer4):

\[6=10x\]Divide both sides by 10\[\frac{6}{10} = \frac{10x}{10}\]\[x=\frac{6}{10}=\frac{3}{5}\](after reducing fraction)